A coil of resistance R, inductance L, and negligib Ans: .354
Answer #1

XL = 2*pi*f*L = 2*pi*60*L

power factor , cos(theta) = R/Z = 0.750

=> impedence , Z = R/0.750 = [R^2 + (XL)^2]^0.5

=> (1.33R)^2 = R^2 + (XL)^2

=> 0.768R^2 = (XL)^2 = (2*pi*60*L)^2

=> L = 2.32R*10^-3

now f = 180 Hz

=> XL = 2*pi*180*L = 180*2*pi*2.32*R*10^-3 = 2.62R

=> impedence . Z =[ R^2 + (XL)^2]^0.5 = [R^2 + (2.62R)^2]^0.5 = 2.80R

=> power factor , cos(theta) = R/Z = R/2.80R = 0.357